OpenStudy (anonymous):

Can anyone help me with my math? Question: Joe tossed 4 into the pot. Their average weight was 6 lb. Then he tossed in 5 more whose average weight was 7 lb. Then he threw in a big one. The overall average weight of all of them was 7 lb. What did the big one weigh?

OpenStudy (anonymous):

(4(6)+5(7)+x)/10=?

OpenStudy (anonymous):

hello???

OpenStudy (anonymous):

Ohh that's how I do it?

OpenStudy (anonymous):

kind of, im trying to help you solve the equation....

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

okay but you need to give me an answer to the equation

OpenStudy (anonymous):

(4(6)+5(7)+x)/10=?

OpenStudy (kagıtucak):

6x4+7x5=49 first sum then add a big one with weight x new sum=49+x now u have 10 pots at the end 49+x/10=7 49+x=70 x=21

OpenStudy (anonymous):

omg. no one every lets people figure it out themselves.

OpenStudy (anonymous):

OH MY GOD I LOVE U GUYS ^_^

OpenStudy (kagıtucak):

but i did a mistake 6x4+7x5=59 not 49 so answer is 11 sorry again

OpenStudy (anonymous):

Its ok ^_^ THank u

OpenStudy (tkhunny):

Tiny hint. NEVER use the symbol "x" to mean multiplication. It is far too confusing. Bigger hint. WRITE DOWN clear definitions of any variable you introduce. Here's a really clear version - just for an example. If "w" is the weight of the large thing that we are tossing into the pot, we have: 4*6 lb + 5*7 lb + w = Total weight of all the things we are tossing into the pot. 10*7 lb = Total weight of all the things we are tossing into the pot. The rest is arithmetic/algebra: 4*6 lb + 5*7 lb + w = 10*7 lb 24 lb + 35 lb + w = 70 lb 59 lb + w = 70 lb w = 70 lb - 59 lb = 11 lb Prove it! (24 + 35 + 11)/10 = (24 + 11 + 35)/10 = (35 + 35)/10 = 70/10 = 7 -- Check!